OpenStudy (anonymous):

3(4x+4)=2(5x+9)-12 Show your work? :3

OpenStudy (snuggielad):

simple and fun...all you have to do is distribute and solve...do you mean=12 though...

OpenStudy (snuggielad):

i got this waynex

OpenStudy (anonymous):

3(4x+4)=2(5x+9)-12 12x + 12 = 10x + 18 -12 12x -10x = 18-12-12 2x = -6 x = -3

OpenStudy (snuggielad):

TOMO DELETE THAT...ITS CHEATING hurry

OpenStudy (snuggielad):

the mods will get on your tail for that...especially with the recent events

OpenStudy (anonymous):

that's not cheating. that's explanation with style

OpenStudy (e.cociuba):

Distribute, Combine like terms, Subtract 12 from both side, subtract 10x from both side, Combine like terms on each side again, And divide both sides by 2 to isolate x. You get x= -3 (If u do it rite)

OpenStudy (e.cociuba):

Thts not cheating.. For some reason u always have problems with the mods... Y?

OpenStudy (snuggielad):

*facepalm* @karatechopper @Preetha please give me a hand

OpenStudy (e.cociuba):

He explained it and gave an answer. Your always dissing on them, no offense but u need to stop/

OpenStudy (snuggielad):

it is cheating...he asked for the answer and show your work...you just gave him an easy way out of learning...you know hes not going to check that...he is just goin to use the answers ans leave

OpenStudy (karatechopper):

.....Its practically helping to me, it shows the steps..

OpenStudy (e.cociuba):

Well than thts his prob not ur's. He was helping.

OpenStudy (snuggielad):

*facepalm*ok..whatever...im not going to argue..i have no say...

OpenStudy (anonymous):

First of all, I'm not a guy. Second of all I HAVE to check my answers on homework or the question won't count.

OpenStudy (e.cociuba):

And srry about calling u a guy @KayKoo2013 My bad :)

OpenStudy (anonymous):

lol it's okay :)

OpenStudy (e.cociuba):

:)